OpenStudy (anonymous):

Which number is between 0.5 and 8/9? Option A: 7/24 Option B: 1/3 Option C: 22/24 Option D: 19/24 This is really weird, because I got the answer of 25/36 so really need help plz?

OpenStudy (anonymous):

D is the correct option

OpenStudy (anonymous):

howd you get that?

OpenStudy (anonymous):

are you allowed to use calculator?

OpenStudy (anonymous):

we cud notice that 8/9=0.888 so any number between 0,5 and 0,8 works .we should express al numbers in decimal form then

OpenStudy (anonymous):

in order to see this

OpenStudy (anonymous):

other wise we cud check using fractions since everything is expressed as frac so \[[1/2\space ,8/9]\] 12/24 is half if any number is more that 12 on the numerator then it is between the 1/2 and 8/9 taking note that 1/3=8/24

OpenStudy (anonymous):

8/9 means he numerator should be less than or equal 21

OpenStudy (anonymous):

since 8/9(24)=21

OpenStudy (anonymous):

so our problem reduces to what numbers are between 12 and 21 7 8 22 19

OpenStudy (anonymous):

obviously we see the result